A car makes a trip due north for three-fourths of the time and due south one-fourth of the time. The average northward velocity has a magnitude of 11 m/s, and the average southward velocity has a magnitude of 21 m/s. What is the average velocity (magnitude and direction) for the entire trip?
Answer #1

let t be the time taken for the entire trip

and y be the dispalcemnt

then average velocity for the entire trip is y/t

then dispalcemnt y = 11*(3t/4) - (21*t/4) = 12*t/4 = 3*t


then required average velocity is y/t = 3 m/s